OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3.

OpenStudy (anonymous):

Point-slope: y – y1 = m(x – x1)

OpenStudy (anonymous):

Plug 6 in for y1 and -1 for x1 and -3 for m, and simplify if required.

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

Does it make sense?

OpenStudy (anonymous):

no can you do it for me ?

OpenStudy (anonymous):

This is the Point-Slope formula y – y1 = m(x – x1) Then all you do is plug it in

OpenStudy (anonymous):

Thats how you get the anwser y - (-1) = -3(x - (-3)) y + 1 = -3(x+3)

OpenStudy (anonymous):

y − 6 = −3(x + 1) is it

OpenStudy (anonymous):

yes like that. D:

OpenStudy (anonymous):

Sorry if my example was confusing.

OpenStudy (anonymous):

thank you